Given the equation of the line: 3x + 7y = 128. When y = 2, what is the value of x?
Answer #1

Given equation, 3x+7y =128

Given y= 2

Substitute y =2 in the given equation

3x+7(2)=128

3x+14=128

3x=128-14

3x =114

x=114/3

x=38

____×____

All the best

Please do support